# NutePrune: Efficient Progressive Pruning with Numerous Teachers for Large Language Models

### v1

@@ Abstract
The considerable size of Large Language Models (LLMs) presents notable deployment challenges, particularly on resource-constrained hardware. Structured pruning, offers an effective means to compress LLMs, thereby reducing storage costs and enhancing inference speed for more efficient utilization. In this work, we study data-efficient and resource-efficient structure pruning methods to obtain smaller yet still powerful models. Knowledge Distillation is well-suited for pruning, as the intact model can serve as an excellent teacher for pruned students. However, it becomes challenging in the context of LLMs due to memory constraints. To address this, we propose an efficient progressive Numerous-teacher pruning method (NutePrune). NutePrune mitigates excessive memory costs by loading only one intact model and integrating it with various masks and LoRA modules, enabling it to seamlessly switch between teacher and student roles. This approach allows us to leverage numerous teachers with varying capacities to progressively guide the pruned model, enhancing overall performance. Extensive experiments across various tasks demonstrate the effectiveness of NutePrune. In LLaMA-7B zero-shot experiments, NutePrune retains 97.17% of the performance of the original model at 20% sparsity and 95.07% at 25% sparsity.
@@
Shengrui Li1 Tsinghua University Microsoft Research Asia lsr22@mails.tsinghua.edu.cn                        Xueting Han2 Microsoft Research Asia chrihan@microsoft.com                        Jing Bai Microsoft Research Asia jbai@microsoft.com
11footnotetext: Work is done during internship at Microsoft Research Asia.22footnotetext: Corresponding author.
@@ 1 Introduction
In this paper, we address the above challenges with an efficient progressive Numerous-teacher pruning method (NutePrune). Our motivation is demonstrated in Figure 1. NutePrune aims to diminish the capacity gap between the full teacher model and the highly sparse student, thereby alleviating the difficulty of distillation Su et al. (2021); Mukherjee et al. (2023); Xiang et al. (2020). Instead of relying solely on a single full teacher, we instruct the student with many teachers with varying sparsity. To achieve this, we formulate pruning as a optimization problem where we learn masks to prune sub-modules while updating model parameters through LoRA Hu et al. (2021). Specially, we load an intact model, serving dual roles as both a teacher and a student. In teacher mode, we incorporate the original model with collected frozen low-sparsity masks and corresponding LoRA modules. And in student mode, we incorporate it with learnable high-sparsity masks and LoRA modules. Since the masks and LoRA modules are highly parameter efficient, we collect and leverage numerous modules with different sparsity to incorporate numerous teachers and progressively prune the student. Our contributions can be summarized as follows:
@@
  • Extensive experiments across perplexity metric, commonsense reasoning, and MMLU demonstrate the effectiveness of our method.
@@ 2 Related Works
For LLMs, SparseGPT Frantar and Alistarh (2023) and WANDA Sun et al. (2023) employ unstructured pruning methods, while N:M sparsity Zhou et al. (2021) is considered semi-structured. Despite the effectiveness of these methods, their intricate structures do not yield significant inference speedup on standard hardware Frantar and Alistarh (2023). Consequently, structured pruning has emerged as a recent consensus. CoFi Xia et al. (2022) and nn pruning Lagunas et al. (2021) are proposed for smaller language models like BERT Devlin et al. (2018), often designed for specific tasks. CoFi loads both the teacher and student models, which is impractical for LLMs. Sheared-LLaMA Xia et al. (2023) proposes pruning LLMs using a dynamic pre-training method, enhancing performance through extensive data and training resources.
@@ 3.1 Efficient Distillation for Structured Pruning
We formulate structure pruning as a constrained optimization problem where we optimize masks and model to prune the structure to a target sparsity while maximizing performance. We use distillation loss instead of original language model loss and use $L_{0}$ regularization to control the pruned sparsity. To mitigate substantial memory consumption during LLM training, we utilize LoRA for model updates, making pruning the process of training these masks and LoRA parameters.
@@
We allow for three types of structured pruning: attention heads, FFN intermediate dimensions, and hidden dimensions. It is achieved by learning corresponding masks $\mathbf{z}_{head},\mathbf{z}_{int},\mathbf{z}_{hid}\in\{0,1\}$. Formally, the multi-head attention module $\operatorname{MHA}(x)$ and feed-forward networks $\operatorname{FFN}(x)$ of layer $l$ are pruned as:
@@
$\begin{split}\operatorname{FFN}^{l}&(X)=\mathbf{z}_{hid}\\ &\cdot W_{D}^{l}\left(\mathbf{z}_{int}^{l}\cdot W_{U}^{l}(X)\cdot\operatorname{Act}(W_{G}^{l}(X))\right)\end{split}$ (2)
where $\operatorname{Att}()$ is the attention module and $\operatorname{Act}()$ is the activation. $W_{D},W_{U},W_{G}$ are down projection, up projection, and gating projection.
@@
All masking variables are learned as real numbers in $[0,1]$ during training. We follow Louizos et al. (2017); Guo et al. (2023) and employ the augmented $L_{0}$ regularization:
$\mathbf{u}$ $\sim U(0,1)$ (4)
$\tilde{\mathbf{s}}$ $=\mathbf{s}\times(r-l)+l$
$\mathbf{z}$ $=\min(1,\max(0,\tilde{\mathbf{s}})),$
@@ 4.1 Experimental Setup
To assess the zero-shot ability of LLMs, we perform zero-shot classification tasks on seven commonsense reasoning benchmarks: BoolQ Clark et al. (2019), PIQA Bisk et al. (2020), HellaSwag Zellers et al. (2019), WinoGrande Sakaguchi et al. (2021), ARC-easy Clark et al. (2018), ARC-challenge Clark et al. (2018), and OpenBookQA (OBQA) Mihaylov et al. (2018). We evaluate the general capcability of LLMs on the perplexity metric with WikiText Merity et al. (2016) and PTB Marcus et al. (1993) dataset. Additionally, We report the results on 5-shot MMLU to evaluate the in-context learning ability Hendrycks et al. (2020), which consists of 57 tasks covering STEM, humanities, social science, etc.
@@
We assess the performance of NutePrune on the LLaMA-1 family Thoppilan et al. (2022), comprising LLaMA-7B, LLaMA-13B. Our experiments primarily center on pruning the LLaMA-7B model to provide a comprehensive comparison with previous studies. We primarily evaluate at sparsity levels of 20% and 50%.
Ratio Tune Method BoolQ PIQA
HellaSwag
WinoGrande
ARC-e
ARC-c
OBQA
@@
0% LLaMA-7B 73.18 78.35
72.99
67.01
67.45
41.38
42.40
63.25 66.39
20% LLM-Pruner 59.39 75.57
65.34
61.33
59.18
37.12
39.80
56.82 59.01
LoRAPrune 57.98 75.11
65.81
59.90
62.14
34.59
39.98
56.50 -
@@
67.96
66.69
63.72
38.05
40.00
59.46 63.03
20%   $\checkmark$ WANDA 65.75 74.70
64.52
59.35
60.65
36.26
39.40
57.23 -
LLM-Pruner 69.54 76.44
68.11
65.11
63.43
37.88
40.00
60.07 61.94
LoRAPrune 65.82 79.31
70.00
62.76
65.87
37.69
39.14
60.05 -
LoRAShear 70.17 76.89
68.69
65.83
64.11
38.77
39.97
60.63 -
Compresso 73.64 75.08
64.77
67.72
66.12
37.54
40.40
60.75 62.60
@@
66.69
64.56
61.49
37.03
39.20
58.71 63.12
NutePrune 68.99 77.20
67.90
65.04
63.76
37.80
40.20
60.13 63.78
50% LLM-Pruner 52.32 59.63
35.64
53.20
33.50
27.22
33.40
42.13 40.94
LoRAPrune 51.78 56.90
36.76
53.80
33.82
26.93
33.10
41.87 -
@@
53.03
57.06
45.45
30.03
36.60
50.35 53.14
50% $\checkmark$ WANDA 50.90 57.38
38.12
55.98
42.68
34.20
38.78
45.43 -
LLM-Pruner 61.47 68.82
47.56
55.09
46.46
28.24
35.20
48.98 48.97
LoRAPrune 61.88 71.53
47.86
55.01
45.13
31.62
34.98
49.71 -
LoRAShear 62.12 71.80
48.01
56.29
47.68
32.26
34.61
50.39 -
Compresso 60.09 66.70
39.31
51.93
48.82
27.82
33.40
46.87 47.43
@@
  • $\star$ includes our reproduced results with the newer version of lm-evaluation-harness that fixes a LLaMA tokenization issue. See Appendix A for detailed results.
Table 1: Zero-shot performance (%) of the compressed LLaMA models. Bold denotes the best average performance at the same setting.
Baselines
Numerous works delve into pruning techniques for LLMs, including SparseGPT Frantar and Alistarh (2023) and WANDA Sun et al. (2023). Considering the benefits of inference acceleration, we mainly focus on structure pruning methods. For open-source methods as far as we know, we implement LLM-Pruner Ma et al. (2023) and Compresso Guo et al. (2023) and conduct detailed comparisons with our NutePrune. For more recent works that are not publicly available, We assess NutePrune using the same settings as theirs for comparison with their reported results. This includes LoRAPrune Zhang et al. (2023) and LoRAShear Chen et al. (2023).
Implementation details
For pruning stage, we sample 20,000 sentences from the C4 Raffel et al. (2020) dataset with a length of 512 tokens. We train with AdamW optimizer, a batch size of 16, and learning rates of 0.1 for masks and 0.001 for LoRA. We prune the model for 7 epochs. We use a linear sparsity schedule to gradually increase the target sparsity from 0 to the target ratio. We set sparsity warmup to 4 epochs for pruning of 20% sparsity and 1 epoch for 50%. The sparsity gap between the teacher and student $g$ is 10% and the snapshot interval $i$ of teachers is 1%. After pruning, we post fine-tune the pruned model on the Alpaca dataset Taori et al. (2023) for 3 epochs. All experiments are conducted on one A100 GPU (80G).
@@ 4.2 Results
Table 1 demonstrates the zero-shot performance on commonsense reasoning tasks for compressed LLaMA models. We mainly present results with the previous version of lm-evaluation-harness for comprehensive comparison, which is widely employed in prior studies Ma et al. (2023). Additionally, we present results with a newer version that addressed a tokenization issue in LLaMA for more accurate evaluation, where we report our reproduced results of publicly available methods. The reported results include experiments for 20%, 25% and 50% sparsity levels, covering scenarios with and without parameter tuning.
The average performance of NutePrune consistently outperforms previous methods across all settings. For pruning without tuning, NutePrune outperforms LLM-Pruner by 2.64%/8.22% at 20%/50% sparsity, underscoring its ability to derive a more effective pruned structure compared to other methods. For pruning with LoRA con-trained, NutePrune improves from 59.46%/50.35% to 61.46%/52.13% at 20%/50% sparsity, indicating co-training with LoRA could help recover model capability damaged by pruning. And with addtional post fine-tuning on Alpaca, notably, it retains 97.17% of the performance of the original model at 20% sparsity and 95.07% at 25% sparsity. NutePrune exhibits more noticeable improvements at higher sparsity levels, proving the effectiveness of our PKD in mitigating the capacity gap between teacher and student. Specifically, NutePrune achieves comparable performance to other methods at 20% sparsity when operating at 25% sparsity. And even without parameter tuning, NutePrune achieves results comparable to other methods at 50% sparsity. These findings suggest that NutePrune is particularly effective in compressing models to higher sparsity levels.
Perplexity
The perplexity (PPL) is a metric reflecting the generation ability of LLMs. The results on WikiText2 and PTB are presented in Table 2. Compresso prunes on an instruction dataset and depends on a meticulously crafted instructing prompt, enabling it to outperform LLM-Pruner on zero-shot tasks at low sparsity but yielding inferior performance on PPL, especially at high sparsity. In contrast, NutePrune utilizes a more general 20k C4 dataset, preserving overall language ability. This distinction allows NutePrune to avoid degradation of generalization and achieve superior performance in terms of PPL compared to both Compresso and LLM-Pruner.
@@
0% LLaMA-7B 5.68 8.81
20% LLM-Pruner 9.96 15.61
[…]
20% $\checkmark$ LLM-Pruner 8.57 12.84
Compresso 10.38 16.14
NutePrune 7.65 12.40
[…]
25% $\checkmark$ NutePrune 7.85 13.10
50% LLM-Pruner 98.1 224.54
[…]
50% $\checkmark$ LLM-Pruner 22.76 33.9
Compresso 59.73 80.38
NutePrune 13.20 21.97
Table 2: Perplexity metric on WikiText2 and PTB.
@@
In addition to evaluating LLaMA-7B, we assess the larger LLaMA-13B with 20% sparsity. As demonstrated in Table 3, our approach yield an average zero-shot performance of 67.51%, which is only 0.12% lower than the full model and 1.75% higher than LLM-Pruner. It also outperforms LLM-Pruner in terms of PPL. Besides, in-context learning is a crucial capability for LLMs Brown et al. (2020), particularly for larger language models such as LLaMA-13B. To assess it, we conduct evaluations using the MMLU with a 5-shot setting. Results demonstrates that NutePrune can also perform well on MMLU metrics with larger models.
Method Zero-Shot (%) WikiText2$\downarrow$ MMLU
0% Baseline 67.63 5.62 0.426
LLM-Pruner 65.76 6.95 0.351
NutePrune 67.51 6.55 0.355
Table 3: Performance of the compressed LLaMA-13B models with 20% sparsity.
@@
We test the inference latency by generating from 64 tokens to 256 tokens on vLLM Kwon et al. (2023), which is a fast and widely deployed library for LLM inference and serving. The results are presented in Table 4. NutePrune achieves latency savings of 11% and 29% at sparsity levels of 20% and 50%, respectively. While LLM-Pruner can save slightly more latency due to its predefined neater structure, it comes at the cost of reduced flexibility in tailoring. However, as sparsity increases, the difference becomes negligible.

### v2

@@ Abstract
The considerable size of Large Language Models (LLMs) presents notable deployment challenges, particularly on resource-constrained hardware. Structured pruning, offers an effective means to compress LLMs, thereby reducing storage costs and enhancing inference speed for more efficient utilization. In this work, we study data-efficient and resource-efficient structure pruning methods to obtain smaller yet still powerful models. Knowledge Distillation is well-suited for pruning, as the intact model can serve as an excellent teacher for pruned students. However, it becomes challenging in the context of LLMs due to memory constraints. To address this, we propose an efficient progressive Numerous-teacher pruning method (NutePrune). NutePrune mitigates excessive memory costs by loading only one intact model and integrating it with various masks and LoRA modules, enabling it to seamlessly switch between teacher and student roles. This approach allows us to leverage numerous teachers with varying capacities to progressively guide the pruned model, enhancing overall performance. Extensive experiments across various tasks demonstrate the effectiveness of NutePrune. In LLaMA-7B zero-shot experiments, NutePrune retains 97.17% of the performance of the original model at 20% sparsity and 95.07% at 25% sparsity. Our code is available at https://github.com/Lucius-lsr/NutePrune.
@@
Shengrui Li1,2†      Junzhe Chen1,2†       Xueting Han2*       Jing Bai2 1Tsinghua University       2Microsoft Research Asia {lsr22, chenjz20}@mails.tsinghua.edu.cn {chrihan, jbai}@microsoft.com
footnotetext: Work done during internship at Microsoft Research Asia.**footnotetext: Corresponding author.
@@ 1 Introduction
Method NutePrune LLM-Pruner KD
GPU Memory (GB) 28.7 35.4 42.1
Table 1: GPU memory consumption during pruning.
In this paper, we address the above challenges with an efficient progressive Numerous-teacher pruning method (NutePrune). Our motivation is demonstrated in Figure 1. NutePrune aims to diminish the capacity gap between the full teacher model and the highly sparse student, thereby alleviating the difficulty of distillation Su et al. (2021); Mukherjee et al. (2023); Xiang et al. (2020). Instead of relying solely on a single full teacher, we instruct the student with many teachers with varying sparsity. To achieve this, we formulate pruning as a optimization problem where we learn masks to prune sub-modules while updating model parameters through LoRA Hu et al. (2021). Specially, we load an intact model, serving dual roles as both a teacher and a student. In teacher mode, we incorporate the original model with collected frozen low-sparsity masks and corresponding LoRA modules. And in student mode, we incorporate it with learnable high-sparsity masks and LoRA modules. Since the masks and LoRA modules are highly parameter efficient, we collect and leverage numerous modules with different sparsity to incorporate numerous teachers and progressively prune the student. And as shown in Table 1, this novel strategy remains highly memory efficient. Our contributions can be summarized as follows:
@@
  • Extensive experiments, including LLaMA-1/2/3 with varying sizes and Mistral, demonstrate the effectiveness of our approach across perplexity, commonsense reasoning, MMLU, and BBH.
@@ 2 Related Works
Pruning Type Speedup No Support No Index
Unstructured $\checkmark$
Semi-Structured $\checkmark$
[…]
Table 2: Structured pruning yield most significant speedup without any special hardware support or additional index storage.
@@
For LLMs, SparseGPT Frantar and Alistarh (2023) and WANDA Sun et al. (2023) employ unstructured pruning methods, while N:M sparsity Zhou et al. (2021) is considered semi-structured. Despite the effectiveness of these methods, their intricate structures do not yield significant inference speedup on standard hardware Frantar and Alistarh (2023) and they need to store additional indexes. As compared in Table 2, structured pruning offers significant advantages, resulting in increased focus on this field in recent works. CoFi Xia et al. (2022) and nn pruning Lagunas et al. (2021) are proposed for smaller language models like BERT Devlin et al. (2018), often designed for specific tasks. CoFi loads both the teacher and student models, which is impractical for LLMs. Sheared-LLaMA Xia et al. (2023) proposes pruning LLMs using a dynamic pre-training method, enhancing performance through extensive data and training resources.
@@ 3.1 Efficient Distillation for Structured Pruning
We formulate structure pruning as a constrained optimization problem where we simultaneously learn masks to prune the structure and update the model to recover ability. To mitigate memory consumption, we utilize LoRA for model updates, making pruning the process of training these masks and LoRA parameters.
@@
Three types of structure are pruned: attention heads, FFN intermediate dimensions, and hidden dimensions. We achieve this by learning masks $\mathbf{z}_{head},\mathbf{z}_{int},\mathbf{z}_{hid}\in\{0,1\}$. Formally, the multi-head attention module $\operatorname{MHA}(x)$ and feed-forward networks $\operatorname{FFN}(x)$ of layer $l$ are pruned as:
@@
$\operatorname{FFN}^{l}(X)=\mathbf{z}_{hid}\cdot W_{D}^{l}\left(\mathbf{z}_{int}^{l}\cdot W_{U}^{l}(X)\cdot W_{G}^{l}(X)\right)$ (2)
where $\operatorname{Att}()$ is the attention module and activation is omitted. $W_{D},W_{U},W_{G}$ are down projection, up projection, and gating projection.
@@
All masking variables are learned as real numbers in $[0,1]$ during training. We follow Louizos et al. (2017); Guo et al. (2023) and employ the augmented $L_{0}$ regularization, which is detailed in Appendix A.
@@ 4.1 Experimental Setup
To assess the zero-shot ability of LLMs, we perform zero-shot classification tasks on seven commonsense reasoning benchmarks: BoolQ Clark et al. (2019), PIQA Bisk et al. (2020), HellaSwag Zellers et al. (2019), WinoGrande Sakaguchi et al. (2021), ARC-easy Clark et al. (2018), ARC-challenge Clark et al. (2018), and OpenBookQA (OBQA) Mihaylov et al. (2018). We evaluate the general capcability of LLMs on the perplexity metric with WikiText Merity et al. (2016). Additionally, to evaluate the in-context learning ability, We report the results on 5-shot MMLU Hendrycks et al. (2020), and 3-shot BBH Suzgun et al. (2022).
@@
NutePrune is applicable across various models of different sizes. We assess the performance of NutePrune on the LLaMA-1 family (7B/13B) Thoppilan et al. (2022), LLaMA-2 Touvron et al. (2023) family (7B/13B), LLaMA-3-8B and Mistral-7B Jiang et al. (2023).
Baselines
Considering the benefits of inference acceleration, we focus on structured pruning. We first replicate conventional methods: Magnitude pruning (MaP) Li et al. (2018), Movement Pruning (MvP) Sanh et al. (2020), and WANDA Sun et al. (2023). For recent open-source methods, we implement LLM-Pruner Ma et al. (2023) and Compresso Guo et al. (2023) and conduct detailed comparisons. For more recent works that are not publicly available, we assess NutePrune using the same settings as theirs. This includes LoRAPrune Zhang et al. (2023) and LoRAShear Chen et al. (2023).
Implementation details
For pruning stage, we sample 20,000 sentences from the C4 Raffel et al. (2020) dataset with a length of 512 tokens. We train with AdamW optimizer, a batch size of 16, and learning rates of 0.1 for masks and 0.001 for LoRA. We prune the model for 7 epochs and a linear sparsity schedule for target sparsity warmup: 4 epochs for 20% sparsity and 1 epoch for 50%. The sparsity gap between the teacher and student $g$ is 10% and the snapshot interval $i$ of teachers is 1%. After pruning, we post fine-tune the pruned model on the Alpaca dataset Taori et al. (2023) for 3 epochs. All experiments are conducted on one A100 GPU (80G).
@@
0% LLaMA-7B 5.68 73.18 78.35 72.99 67.01 67.45 41.38 42.40 63.25 66.39
20% LLM-Pruner 9.96 59.39 75.57 65.34 61.33 59.18 37.12 39.80 56.82 59.01
LoRAPrune - 57.98 75.11 65.81 59.90 62.14 34.59 39.98 56.50 -
@@
20% Tuned MaP 12.67 60.00 76.12 65.43 60.93 60.31 37.80 39.80 57.20 60.05
MvP 10.52 64.50 73.50 62.50 61.80 62.42 36.95 37.80 57.07 58.76
WANDA - 65.75 74.70 64.52 59.35 60.65 36.26 39.40 57.23 -
LLM-Pruner 8.57 69.54 76.44 68.11 65.11 63.43 37.88 40.00 60.07 61.94
LoRAPrune - 65.82 79.31 70.00 62.76 65.87 37.69 39.14 60.05 -
LoRAShear - 70.17 76.89 68.69 65.83 64.11 38.77 39.97 60.63 -
Compresso 10.38 73.64 75.08 64.77 67.72 66.12 37.54 40.40 60.75 62.60
@@
NutePrune 7.85 68.99 77.20 67.90 65.04 63.76 37.80 40.20 60.13 63.78
50% LLM-Pruner 98.10 52.32 59.63 35.64 53.20 33.50 27.22 33.40 42.13 40.94
LoRAPrune - 51.78 56.90 36.76 53.80 33.82 26.93 33.10 41.87 -
@@
50% Tuned MaP 33.18 39.69 66.81 42.49 50.67 49.32 30.63 31.40 44.43 46.33
MvP 27.62 59.94 63.06 40.98 55.64 44.07 26.79 31.80 46.04 46.23
WANDA - 50.90 57.38 38.12 55.98 42.68 34.20 38.78 45.43 -
LLM-Pruner 22.76 61.47 68.82 47.56 55.09 46.46 28.24 35.20 48.98 48.97
LoRAPrune - 61.88 71.53 47.86 55.01 45.13 31.62 34.98 49.71 -
LoRAShear - 62.12 71.80 48.01 56.29 47.68 32.26 34.61 50.39 -
Compresso 59.73 60.09 66.70 39.31 51.93 48.82 27.82 33.40 46.87 47.43
@@
  • $\star$ includes results with the newer version of lm-evaluation-harness. See Appendix B for detail.
Table 3: Performance (%) of the compressed LLaMA-7B models.
[…]
0% LLaMA-2-7B 5.47 77.74 79.11 75.97 69.06 76.35 46.33 44.20 66.97
20% LLM-Pruner 12.94 50.55 75.46 67.18 65.67 67.38 38.14 38.40 57.54
NutePrune 8.74 77.06 76.66 70.56 65.59 71.97 42.58 42.40 63.83
50% LLM-Pruner 24.47 54.13 68.06 46.71 51.54 50.97 25.85 34.00 47.30
NutePrune 12.94 66.24 70.83 57.04 59.51 58.46 31.97 34.00 54.01
Table 4: Performance (%) of the compressed LLaMA-2-7B models.
[…]
0% Mistral-7B 5.25 83.73 82.26 81.05 74.19 80.89 53.84 43.80 71.39
20% LLM-Pruner 7.50 77.52 78.13 73.64 69.46 72.59 46.41 41.80 65.65
NutePrune 7.06 78.29 80.41 75.57 68.82 76.35 45.05 42.20 66.67
50% LLM-Pruner 30.51 62.48 66.59 48.00 56.51 52.61 28.07 29.80 49.15
NutePrune 12.29 63.64 72.63 57.95 61.25 62.46 35.41 33.80 55.31
Table 5: Performance (%) of the compressed Mistral-7B models.
@@ 4.2 Results
0% LLaMA-3-8B 6.14 81.04 80.85 79.18 73.40 80.13 53.16 44.60 70.34
20% LLM-Pruner 10.06 71.50 77.97 70.49 68.75 72.35 42.83 38.40 63.18
NutePrune 9.51 78.65 79.76 73.74 70.09 76.22 45.90 43.40 66.82
50% LLM-Pruner 27.37 41.59 67.46 46.53 55.64 49.71 27.22 31.60 45.68
NutePrune 21.72 65.20 68.72 52.30 58.64 53.41 29.44 35.00 51.82
Table 6: Performance (%) of the compressed LLaMA-3-8B models.
Ratio Method Avg. Zero-Shot (%) WikiText2$\downarrow$ MMLU (5-shot) BBH (3-shot)
0% LLaMA-13B 67.63 5.62 46.90 37.72
20% LLM-Pruner 65.76 6.95 29.57 15.77
NutePrune 67.51 6.55 39.37 31.99
0% LLaMA-2-13B 68.80 5.30 54.86 39.53
20% LLM-Pruner 65.87 7.25 31.26 25.20
NutePrune 67.39 6.86 45.78 31.22
Table 7: Performance of the compressed LLaMA-13B and LLaMA-2-13B models with 20% sparsity.
Table 3 demonstrates PPL and zero-shot performances on commonsense reasoning tasks for compressed LLaMA-7B models. The reported results include experiments for 20%, 25% and 50% sparsity levels, covering scenarios with and without parameter tuning.
The average performance of NutePrune consistently outperforms previous methods across all settings. For pruning without tuning, NutePrune outperforms LLM-Pruner by 2.64%/8.22% at 20%/50% sparsity, underscoring its ability to derive a more effective pruned structure compared to other methods. For pruning with LoRA con-trained, NutePrune improves from 59.46%/50.35% to 61.46%/52.13% at 20%/50% sparsity, indicating co-training with LoRA could help recover model capability damaged by pruning. And with additional post fine-tuning on Alpaca, notably, it retains 97.17% of the performance of the original model at 20% sparsity and 95.07% at 25% sparsity.
Table 4, 5 and 6 further demonstrates performances for compressed LLaMA-2-7B, Mistral-7B and LLaMA-3 models. At the same sparsity, multi-query attention models experience a more significant performance decline. Nevertheless, NutePrune consistently outperforms LLM-Pruner, proving our method is applicable across various models. Noticeable improvements are observed at higher sparsity levels, proving the effectiveness of our PKD in mitigating the capacity gap during distillation.
@@
We assess larger models: LLaMA-13B and LLaMA-2-13B with 20% sparsity. To evaluate the ability of these stronger models, we further assess their in-context learning ability with MMLU and BBH Brown et al. (2020). As demonstrated in Table 7, our approach yield an average zero-shot commonsense reasoning performance of 67.51% and 67.39%, which is only slightly lower than the full model and much higher than LLM-Pruner. It also outperforms LLM-Pruner in terms of PPL in WikiText2. For in-context learning ability, NutePrune achieves a score of 36.37 MMLU and 31.99 BBH in LLaMA-13B, and 45.78 MMLU and 31.22 BBH in LLaMA-2-13B. The slight decline in performance compared to the full model is acceptable, indicating that NutePrune maintains sufficient in-context learning capability. Additionally, when compared to LLM-Pruner, our advantages are clearly evident.
@@
We test the inference latency by generating from 64 tokens to 256 tokens on vLLM Kwon et al. (2023), which is a fast and widely deployed library for LLM inference and serving. The results are presented in Table 8. NutePrune achieves latency savings of 11% and 29% at sparsity levels of 20% and 50%. While LLM-Pruner save slightly more latency due to its predefined neater structure, it comes at the cost of reduced flexibility in tailoring. As sparsity increases, the difference becomes negligible.